OpenStudy (anonymous):

Let A be a square matrix and I is the unit matrix i) Assume A^2 = 0. Show that (I - A)^1 = I + A. ii ) Assume A^3 = 0. Show that (E - A)^-1 = E + A + A^2 . iii ) Determine the inverse of the matrix 1 2 -1 0 1 3 0 0 1 using one of the above equations . iv ) Suppose A ^ k = 0 , k ≥ 2. Determine an expression for (E - A) ^ - 1

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino is i) something like this \[(I-A)^{-1}=I+A \leftarrow \rightarrow (I-A)*(I-A)^{-1}=(I-A)*(I+A)\] \[(I-A)*(I-A)^{-1} = I\] \[I=I^2+I*A-A*I-A^2\] since A^2=0 and I^2=I we can neglect that, so we have: \[I = I + I*A-A*I\]

OpenStudy (anonymous):

\[I-I=A(I-I) <=> 0 = 0\]

OpenStudy (anonymous):

sorry on ii) it should be ii ) Assume A^3 = 0. Show that (I - A)^-1 = I + A + A^2 and NOT ii ) Assume A^3 = 0. Show that (E - A)^-1 = E + A + A^2

OpenStudy (michele_laino):

we can write this: \[\Large \begin{gathered} \left( {I - A} \right){\left( {I - A} \right)^{ - 1}} = \left( {I - A} \right)\left( {I + A} \right) = \hfill \\ \hfill \\ = {I^2} + IA - IA - {A^2} = I - A + A + 0 = I \hfill \\ \end{gathered} \] and: \[\Large \begin{gathered} {\left( {I - A} \right)^{ - 1}}\left( {I - A} \right) = \left( {I + A} \right)\left( {I - A} \right) = \hfill \\ \hfill \\ = {I^2} - IA + IA - {A^2} = I - A + A + 0 = I \hfill \\ \end{gathered} \] and the proof is complete

OpenStudy (michele_laino):

I have used this property: \(IA=AI\)

OpenStudy (zarkon):

why are you using \((I - A)^{-1} = I + A.\) in the proof?

OpenStudy (anonymous):

who?

OpenStudy (michele_laino):

since the question asks to check that if this Matrix: \(I+A\) is the inverse of \(I-A\)

OpenStudy (michele_laino):

and I showed that such statement is true

OpenStudy (zarkon):

you only know that \( A^2 = 0\)...you are trying to show that \((I - A)^{-1} = I + A\)

OpenStudy (michele_laino):

yes! Nevertheless you can use the thesis since the exercise asks for a check

OpenStudy (zarkon):

start with \[(I-A)(I+A)=\cdots=I\] then multiply both sides by \[(I-A)^{-1}\]

OpenStudy (anonymous):

Is it okey or not do what I just did? I think I just did the same as Michele, right? but what to do on iii)?

OpenStudy (zarkon):

I'll give you a hint...use part (ii) to do part (iii)

OpenStudy (anonymous):

so @Zarkon, If I use (I-A)^-1 = E+A and the matrix as A in the equation \[A=\left[\begin{matrix}1 & 2 &-1\\ 0 & 1&3\\ 0 & 0&1\end{matrix}\right]\] and \[I=\left[\begin{matrix}1 & 0 &0\\ 0 & 1&0\\ 0 & 0&1\end{matrix}\right]\] \[(\left[\begin{matrix}1 & 0 &0\\ 0 & 1&0\\ 0 & 0&1\end{matrix}\right] - \left[\begin{matrix}1 & 2 &-1\\ 0 & 1&3\\ 0 & 0&1\end{matrix}\right])^{-1}=\left[\begin{matrix}1 & 0 &0\\ 0 & 1&0\\ 0 & 0&1\end{matrix}\right] + \left[\begin{matrix}1 & 2 &-1\\ 0 & 1&3\\ 0 & 0&1\end{matrix}\right]\] \[\left[\begin{matrix}0 & -2 &1\\ 0 & 0&-3\\ 0 & 0&0\end{matrix}\right]^{-1}=\left[\begin{matrix}2 & 2 &-1\\ 0 & 2&3\\ 0 & 0&2\end{matrix}\right]\]

OpenStudy (anonymous):

@Michele_Laino check my iii)

OpenStudy (anonymous):

What I wanna do next is to multiply the left hand side by the not inverted one and the same to the right hand side, but thats not the inversion

OpenStudy (anonymous):

or am I doing completely wrong?

OpenStudy (anonymous):

the inverted A should be 1 -2 7 0 1 -3 0 0 1

OpenStudy (michele_laino):

I checked that \[\huge {A^3} \ne O\]

OpenStudy (michele_laino):

furthermore: \[\huge {A^2} \ne O\]

OpenStudy (anonymous):

what?

OpenStudy (michele_laino):

if: \[A = \left( {\begin{array}{*{20}{c}} 1&2&{ - 1} \\ 0&1&3 \\ 0&0&1 \end{array}} \right) \Rightarrow {A^3} \ne O,\;{A^2} \ne O\]

OpenStudy (zarkon):

Let \[A=\left[\begin{array}{ccc} 0 & 2 & -1 \\ 0 & 0 & 3 \\ 0 & 0 & 0 \\ \end{array}\right]\] then \[1+A=\left[\begin{array}{ccc} 1 & 2 & -1 \\ 0 & 1 & 3 \\ 0 & 0 & 1 \\ \end{array}\right]\]

OpenStudy (zarkon):

Then \[A^3=0\]

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

So I have to add 1 to both left and right hand side?

OpenStudy (zarkon):

let \[A=-\left[\begin{array}{ccc} 0 & 2 & -1 \\ 0 & 0 & 3 \\ 0 & 0 & 0 \\ \end{array}\right]\] that will work better

OpenStudy (michele_laino):

I write this: \[\begin{gathered} \left( {\begin{array}{*{20}{c}} 1&2&{ - 1} \\ 0&1&3 \\ 0&0&1 \end{array}} \right) = I - A \Rightarrow \hfill \\ \hfill \\ \Rightarrow A = I - \left( {\begin{array}{*{20}{c}} 1&2&{ - 1} \\ 0&1&3 \\ 0&0&1 \end{array}} \right) \hfill \\ \end{gathered} \]

OpenStudy (zarkon):

then I-A is your matrix and its inverse is \[I+A+A^2\]

OpenStudy (zarkon):

you can do it they I originally said but then the inverse is \[I-A+A^2\]

OpenStudy (anonymous):

But I cant just pick a matrix? i am given one in the task?

OpenStudy (michele_laino):

then the inverse matrix, is: \[{\left( {\begin{array}{*{20}{c}} 1&2&{ - 1} \\ 0&1&3 \\ 0&0&1 \end{array}} \right)^{ - 1}} = I + A = I + I - \left( {\begin{array}{*{20}{c}} 1&2&{ - 1} \\ 0&1&3 \\ 0&0&1 \end{array}} \right)\]

OpenStudy (anonymous):

iii ) Determine the inverse of the matrix 1 2 -1 0 1 3 0 0 1 using one of the above equations .

OpenStudy (anonymous):

1 -2 7 0 1 -3 0 0 1

OpenStudy (michele_laino):

yes! I have applied the statement of part ii

OpenStudy (michele_laino):

sorry my Internet Explorer 11, has crashed

OpenStudy (anonymous):

could you explain a little on IV?

OpenStudy (michele_laino):

yor inverse matrix is correct!

OpenStudy (michele_laino):

for part iv, we have to apply the Mathematical induction principle

OpenStudy (michele_laino):

we can say that, if \(A^k=0\) then: \[\Large {\left( {I - A} \right)^{ - 1}} = I + A + ... + {A^{k - 1}}\] proof: we have this computation: \[\begin{gathered} {\left( {I - A} \right)^{ - 1}} = I + A + ... + {A^{k - 1}} \hfill \\ \hfill \\ \left( {I - A} \right)\left( {I + A + ... + {A^{k - 1}}} \right) = \hfill \\ \hfill \\ = I + A + ... + {A^{k - 2}} + {A^{k - 1}} + \hfill \\ \hfill \\ - A - {A^2} - ... - {A^{k - 1}} - {A^k} = \hfill \\ \hfill \\ = I - {A^k} = I \hfill \\ \end{gathered} \] similarly for this part: \[{\left( {I - A} \right)^{ - 1}}\left( {I - A} \right) = \left( {I + A + ... + {A^{k - 1}}} \right)\left( {I - A} \right)\]